OpenStudy (ria23):

Can someone check this quiz for me? 1.) If f(x) = 8x and g(x) = 2x + 1, find (fog)(x) A.) 16x^2 + 8x B.) 16x + 1 C.) 16x + 8 <-- (my answer) D.) 10x + 1 2.) If h(x) = (fog)(x) and h(x) = 5(x+1)^3, find f(x) and g(x) A.) f(x) = x+1; g(x) = 5x^3 B.) f(x) = 5x^3; g(x) = x+1<-- (my answer) C.) f(x) = (x + 1)^3; g(x) = x + 1 D.) f(x) = 5x^3; g(x) = (x + 1)^3 ~Will post the rest in the comments~

OpenStudy (anonymous):

1) is right

OpenStudy (anonymous):

2) is right also

OpenStudy (ria23):

3.) for f(x) = 4x + 1 and g(x) = x^2 - 5, find (f + g)(x) A.) x^2 + 4x + 6 B.) 4x^2 - 19 c.) 4x^3 - 4 D.) x^2 + 4x - 4 <-- (my answer) 4.) ~I did something wrong on this one. ;-;~ For f(x) = 2x + 1 and g(x) = x^2 - 7, find (f - g)(x) A.) x^2 - 2x - 8 B.) 2x^2 - 15 C.) 8 + 2x - 2x D.) -x^2 + 2x - 6 (I got x^2 + 2x +8)

OpenStudy (ria23):

@mathstudent55 Would yhu mind checking these for me real quick. ;c

OpenStudy (ria23):

@mathmale